OpenStudy (anonymous):

Mickey and Minnie Mouse can eat a piece of cheese in 9 minutes. Mickey alone can eat the cheese in 16 minutes. How long would it take Minnie alone to eat the cheese?

OpenStudy (anonymous):

1/16 + 1/x = 1/9

OpenStudy (anonymous):

adding rational expressions ? thanks

OpenStudy (anonymous):

9x + 144 = 16x 7x = 144 x = 20.43 minutes

OpenStudy (anonymous):

yw